OpenStudy (rational):

[UNSOLVED] Prove/Disprove : No term in the below infinite sequence is a prime \(\large 343,~ 34343, ~3434343, ~343434343, \cdots \)

OpenStudy (rational):

Here is what I have tried so far prime factorization of first few terms shows some pattern in \(\mod 3\) : \[ \large \begin{array}\\ \color{Red}{343} &\color{Red}{: 7^3}\\ 34343 &: 61\times 563\\ \color{green}{3434343} &\color{green}{: 3\times 11^2\times 9461}\\ \color{red}{343434343} &\color{Red}{: 7\times 521\times 94169}\\ 34343434343 &: 47\times 79\times 9249511\\ \color{green}{3434343434343} &\color{green}{: 3^2\times 19\times 29\times 67\times10336531}\\ \color{red}{343434343434343} &\color{Red}{: 7\times 151\times 324914232199}\\ 34343434343434343 &: 5638147\times 6091262669 \\ \end{array} \] However the \(\large a_{3k+1}\) terms have no apparent pattern yet

OpenStudy (solomonzelman):

you write "43" whatever number of times in front of 343.

OpenStudy (kainui):

Perhaps you should consider doing the same thing you did, but look for patterns among the different adding of just the sequence "434343" instead of just "43" since I imagine there might be a pattern within these starting with 34343. Is this a programming class? Perhaps you can come across a counter example and just disprove it and save some real effort haha.

OpenStudy (kainui):

I noticed just now that you can always express these numbers as the sum of two number multiplied, that might make it simpler: \[\LARGE 343 = 3(101)+4(10) \\ \LARGE 34343 = 3(10101)+4(1010) \\ \LARGE 3434343 = 3(1010101)+4(101010) \]

OpenStudy (rational):

thats exactly how it can be concluded that the terms \(\large a_{3k}, a_{3k+2}\) \(k\ge 0\) are composite : for \(\mod 3\), we consider factors of \(434343\) : 3x7x13x37x43 since \(\large a_{3k+3} = 10^6a_{3k} + 434343\) , if we can show that \(\large a_{3k}, a_{3k+1}, a_{3k+2}\) are divisible by any one of the factors of \(434343\), then we're done. but only two terms have common factors : {3, 7} . \(\large a_{3k+1}\) terms are diverging forever without any sharing any factors with the \(43434343... \) sequence

OpenStudy (rational):

the \(\large a_{3k+1}\) terms i am talking about in previous reply are in thick below : \[\large \begin{array}\\ \color{gray}{343} &\color{gray}{: 7^3}\\ 34343 &: 61\times 563\\ \color{gray}{3434343} &\color{gray}{: 3\times 11^2\times 9461}\\ \color{gray}{343434343} &\color{gray}{: 7\times 521\times 94169}\\ 34343434343 &: 47\times 79\times 9249511\\ \color{gray}{3434343434343} &\color{gray}{: 3^2\times 19\times 29\times 67\times10336531}\\ \color{gray}{343434343434343} &\color{gray}{: 7\times 151\times 324914232199}\\ 34343434343434343 &: 5638147\times 6091262669 \\ \end{array}\] the terms in gray are divisible by either 3 or 7, so we're done with them.

OpenStudy (kainui):

Yes, that's what I am talking about in my very first response.

OpenStudy (rational):

yes i get that :) I am not even sure if this is the right approach for handling those \(\large a_{3k+1}\) terms they look very hard >.< btw, i have tested all the "434343.." sequence upto 30 digits for common factors. No success :(

OpenStudy (kainui):

Ahhh, well I'm trying more stuff. I'm trying to see if I can link the 3k+1 terms to the ones that are multiples of 3 or 7 somehow so that we can get those by some sort of extension. Hmm...

OpenStudy (kainui):

lol @ikram002p

OpenStudy (ikram002p):

ikr , this question drive me crazy within these days

OpenStudy (kainui):

At first when I saw this I thought it was just like the same as that other question of yours ikram, showing that 1111, 111111, etc... are all not prime.

OpenStudy (mathmate):

:) (just to keep track of this thread!)

OpenStudy (anonymous):

erm about to reveal :P

OpenStudy (kainui):

I want to try to solve this again!

OpenStudy (anonymous):

@ganeshie8 xD we can right then as difference of squares \( x^2-y^2\) s.t \(x-y\neq1\) and x,y are positive natural

OpenStudy (anonymous):

them* not then

ganeshie8 (ganeshie8):

yes every odd integer can be written as difference of squares, how is it going to help ?

OpenStudy (anonymous):

well , any odd can be written but its prime when x-y=1 , right ? then we need to find such relation to show x-y \neq 1 means its not prime :D

OpenStudy (anonymous):

note that we have all of them are 3 mod 4

OpenStudy (anonymous):

so we would have 4n+3=(4n+3)^2-(20n+6) we need some trick to show for 20n+6 is perfect square for values of n we have hmm

OpenStudy (anonymous):

hmm seems mod 4 dont work , i'll try if this gonna work for some mod :|

OpenStudy (mathmate):

@Marki (just a digression) when (x-y)=1, (x-y)(x+y) does not have to be prime, e.g. (5-4)(5+4)=9 but (x-y) ne 1 does imply (x-y)(x+y) is composite. :)

OpenStudy (anonymous):

ok ok but the opposite is true when x-y is not 1 then its not prime , right ?

OpenStudy (mathmate):

Absolutely. The above remark does not change any of your subsequent arguments. In fact the contrapositive is when (x-y)(x+y) is prime, (x-y)=1 is probably what you meant.

OpenStudy (anonymous):

for 9 try x=3 y=0 instead

OpenStudy (anonymous):

yes this is what im standing for :) showing its not prime not the inverse

OpenStudy (anonymous):

wait i made mistake it should be 4n+3=(4n+3)^2-(16n^2+20n+6)

ganeshie8 (ganeshie8):

you got \(a_n \equiv 3 \pmod 4\) because the last two digits of every term in the sequence is \(43\) and \(43\equiv 3 \pmod 4\) ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

how did you get 4n+3=(4n+3)^2-(16n^2+20n+6) ?

OpenStudy (anonymous):

( 4n+3)^2=16n^2+24 n+9 (4n+3)^2-(4n+3)=16n^2+20n+6

ganeshie8 (ganeshie8):

oh just algebra.. fine, im wid u

OpenStudy (anonymous):

but im not sure if this work , its just an idea

ganeshie8 (ganeshie8):

it wont work because there are infinitely many primes of form 4n+3 so its not possible to factor 4n+3 into two non trivial factors always

OpenStudy (anonymous):

no i dont wanna prove in general as i said for n i have only

OpenStudy (anonymous):

i would have this sequence as n (note that numbers in black ) 8585 8585858585 8585858585858585 8585858585858585858585 ......

ganeshie8 (ganeshie8):

ohk you're going to relate \(n\) with the terms in our sequence, makes sense..

OpenStudy (anonymous):

xD but this one failed from first term im trying to find some other pattern like this hoping :| it works

OpenStudy (anonymous):

hmm it dint work for any mod up to 26 i only tried 34343

OpenStudy (anonymous):

i think its dumb way to think of any way :|

ganeshie8 (ganeshie8):

im not even getting ideas, you're atleast trying something !

OpenStudy (anonymous):

i'll try mod something odd

OpenStudy (kainui):

Another form of representation might help, but it also might not. \[\Large -4 *10^{2n+1}+43 \frac{10^{2n+2}-1}{10^2-1}\]

OpenStudy (kainui):

Playing around and letting x=10^2n I discovered kind of a pattern, but it might not be anything special: \[\Large \frac{1}{99} \left( -4x 10^3 +43x10^2+4x10^1-43* 10^0\right)\] Notice the last term doesn't contain x (or 10 really) but it kind of has this alternating thing going on with the signs and values like maybe it's the solution of some kind of polynomial or something... I don't know.

OpenStudy (anonymous):

do u still need help?

OpenStudy (mathmath333):

\(\)